HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

HIGH-RESOLUTION STUDY OF THE ROTATIONAL SPECTRUM OF 2-AMINOPROP-2-ENENITRILE

\underline{K. LUKOVÁ}, Institute of Physics, University of Kassel, Heinrich-Plett-Strasse 40 D-34132 Kassel, Germany; and Department of Analytical Chemistry, University of

Chemistry and Technology, Technická 5, 166 28, Prague, Czech Republic; L. KOLESNIKOVÁ, J. KOUCKÝ, P. KANIA, Š. URBAN, Department of

Analytical Chemistry, University of Chemistry and Technology, Technická 5, 166 28, Prague, Czech Republic; J.C. GUILLEMIN, Institut des Sciences Chimiques de Rennes, CNRS UMR 6226, Université Rennes 1, 263 Avenue Général Leclerc,

35700 Rennes, France

The high-resolution rotational spectrum of 2-aminopropen-2-enitrile (APN) was recorded in two frequency ranges: 204–240 and 292–328 GHz employing the Prague semiconductor millimetre-wave spectrometer^{1}. Several hundreds of a- and b-type transitions corresponding to the 0^{+} and 0^{−} inversion states arising from the inversion motion of the amino group were identified and assigned. Bauder et al.^{2} estimated the energy difference between the energy levels to be 54 ± 22 cm^{−1} based on the relative intensity measurements. Since no c-type transitions connecting the inversion states were assigned in the present spectrum, the value of ΔE separating the levels could not be derived with a better accuracy. Hence, the rotational transitions from both vibrational states were analysed separately by means of the A-reduced Hamiltonian using the SPCAT/SPFIT program package^{3}. The experimentally determined molecular parameters of APN are presented in comparison with their quantum chemical counterparts.

1doi:10.1016/j.molstruc.2006.02.025, P. Kania, L. Stříteská, M. Šimečková, Š. Urban, J. Mol.

Struct., 795, 209-218, (2006).

2doi:10.1016/0009-2614(83)85003-9, A. Bauder, T.-K. Ha, Chem. Phys. Lett., 97, 135-140, (1983).

3doi:10.1016/0022-2852(91)90393-O, H. M. Pickett, J. Mol. Spectrosc., 148, 371-377, (1991).

p-number: p209  Submitted on Thu, 15 Jun 2023 22:44:35 +0200